OpenStudy (anonymous):

Factor. Show Your Work x^2 - 12x + 27

OpenStudy (amistre64):

-9(-3) = 27 -9-3 = -12 (x-3)(x-9)

OpenStudy (anonymous):

thanx love

OpenStudy (amistre64):

hope your learning a little something along the way ;)

OpenStudy (anonymous):

yup yup i appreciate it